# Efficient Best-of-Both-Worlds Algorithms for Contextual Combinatorial Semi-Bandits

### v1

@@ A.2 Proofs and Auxiliary Results for Section 3
    $f(x):[0,1]\rightarrow\mathbb{R}$ is defined through $f(x)=x\log x-x$ for $x\in(0,1]$ and $f(0)=0$, which is continuous on $[0,1]$ and differentiable on $(0,1]$. Note that although $\log x$ is not defined at $x=0$, we have $\lim_{x\rightarrow 0^{+}}x\log x=0$. Observe that its first derivative for $x\in(0,1]$ is $\ln x$. It then follows that $(f^{\prime})^{-1}(-\lambda-c_{k})=e^{-\lambda-c_{k}}$ for all $k\in[K]$. In addition, note that $z=f^{\prime}(x)=\ln x$ ranges over the compact interval $z\in(-\infty,0]$. On that interval, the derivative of the inverse $\partial e^{z}/\partial z=e^{z}$ satisfies $e^{z}\leq e^{0}=1$ for all $z\in(-\infty,0]$. By the mean-value theorem, for any $z_{1},z_{2}\in[-\infty,0]$ there exists $c$ between them such that $\left|e^{z_{1}}-e^{z_{2}}\right|=e^{c}|z_{1}-z_{2}|\leq 1|z_{1}-z_{2}|$. Hence, the function $(f^{\prime})^{-1}(-\lambda-c_{k})=e^{-\lambda-c_{k}}$ is $1$-Lipschitz.
@@
    $f(x)=-\sqrt{x}$ on $(0,1]$. Similar to previous calculations, we have $x=(f^{\prime})^{-1}(z)=1/(4z^{2})$. Note that it must hold $z<0$ for the expression to be defined properly. Hence, with $z=-\lambda-c_{k}<0$, for all $k\in[K]$, we have $(f^{\prime})^{-1}(-\lambda-c_{k})=1/(4(-\lambda-c_{k})^{2})$. As $x\in(0,1]$ we have $z=-1/(2\sqrt{x})\in(-\infty,-1/2]$. On that interval, the derivative of the inverse is:
@@
    On the interval $z\in\left(-\infty,-\frac{1}{2}\right]$, the largest magnitude of derivative of the inverse map $|(f^{\prime})^{-1}(z)|$ occurs at the smallest $|z|$, namely at $|z|=1/2$. This follows from the fact that $|(f^{\prime})^{-1}(z)|$ is decreasing as $|z|$ grows. Then, $\sup|(f^{\prime})^{-1}(z)|=4$, on $z\in(-\infty,-1/2]$. So, $|(f^{\prime})^{-1}(z_{1})-(f^{\prime})^{-1}(z_{2})|\leq 4|z_{1}-z_{2}|$, for all $z_{1},z_{2}\in(-\infty,-1/2]$ Therefore, $(f^{\prime})^{-1}$ is $4$-Lipschitz.
@@
$f^{\prime}(k/K)\leq-\underline{\lambda}-c_{k}\quad\forall k\in[K].$ (11)
As $f$ is strictly convex, its first derivative $f^{\prime}$ is strictly increasing, which in turn implies that its inverse $(f^{\prime})^{-1}$ is also strictly increasing. Applying the inverse function to both sides of (11) yields
$(f^{\prime})^{-1}(-\underline{\lambda}-c_{k})\geq(f^{\prime})^{-1}\left(f^{\prime}\left(\frac{k}{K}\right)\right)=\frac{k}{K}\quad\forall k\in[K].$
Summing over $k=1,\dots,K$ gives $\sum_{k=1}^{K}(f^{\prime})^{-1}(-\underline{\lambda}-c_{k})\geq k$. A similar argument applied to $\bar{\lambda}$ gives $\sum_{k=1}^{K}(f^{\prime})^{-1}(-\bar{\lambda}-c_{k})\leq k$. Thus, by the intermediate value theorem, there exists at least one $\lambda^{*}\in[\underline{\lambda},\bar{\lambda}]$ such that
$\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})=0.$
@@
where $\Delta=k-\sum_{k=1}^{K}(f^{\prime})^{-1}(-c_{k}-\lambda^{\star})$. For each coordinate $k$, the error induced by the difference between $\lambda$ and $\lambda^{\star}$ is dominated by the Lipschitz property (12) through
@@
which in turn implies that $(f^{\prime})^{-1}(-\lambda-c_{k})\geq(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})-\varepsilon/(2\sqrt{K})$. Summing over $k=1,\dots,d$ and by defining $\Delta_{\lambda}=k-\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda-c_{k})$ yields:
$\Delta_{\lambda}\leq k-\left[\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})-\frac{\varepsilon\sqrt{K}}{2}\right].$
By the definition of $\Delta$ we can bound $\Delta_{\lambda}$ as $\Delta_{\lambda}\leq\Delta+\varepsilon\sqrt{K}/2$. In the worst-case scenario when the computed error $\Delta_{\lambda}$ is close to 0, the inequality above reduces to $\Delta\geq\Delta_{\lambda}-\varepsilon\sqrt{K}/2$. Since $\Delta_{\lambda}$ is nearly 0 or in the worst-case non-positive, this yields the bound $\Delta\geq-\varepsilon\sqrt{K}/2$. Therefore, from (13), the error in each coordinate satisfies $|a_{k}-a_{k}^{\star}|\leq\varepsilon/\sqrt{K}$, and therefore $\|a-a^{\star}\|_{2}\leq\varepsilon$. Thus, the claim follows. ∎
@@
Define the exact residual $g(\lambda)=k-\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda-c_{k})$ and the approximate residual $\tilde{g}(\lambda)=k-\sum_{k=1}^{K}\tilde{y}_{k}$, where each $\tilde{y}_{k}$ satisfies $|\tilde{y}_{k}-(f^{\prime})^{-1}(-\lambda-c_{k})|\leq\tau$. To measure the oracle’s perturbation, set $E(\lambda)=\tilde{g}(\lambda)-g(\lambda)=-\sum_{k=1}^{K}\left[\tilde{y}_{k}-(f^{\prime})^{-1}(-\lambda-c_{k})\right]$. Since this is a sum of $K$ individual errors, the triangle inequality gives $|E(\lambda)|\leq\sum_{k=1}^{K}|\tilde{y}_{k}-(f^{\prime})^{-1}(-\lambda-c_{k})|\leq K\tau$.
@@
Provided that $\min\{g(\underline{\lambda}),\,-g(\bar{\lambda})\}>K\tau$, the sign test on $\tilde{g}$ still selects the correct subinterval; hence the bisection converges with each evaluation of $g$ perturbed by at most $\pm K\tau$. After $l$ iterations, the midpoint $\lambda_{l}$ satisfies $\lvert\lambda_{l}-\lambda^{\star}\rvert\;\leq\;\Delta_{0}/2^{l+1}$, with $\Delta_{0}=\bar{\lambda}-\underline{\lambda}$. Upon termination at $\lambda_{l}$, define $\tilde{a}_{k}=\tilde{y}_{i}(\lambda_{l})$ for all $k\in[K]$. Finally, subtract the mean of $\tilde{a}$ to ensure that $\sum_{k=1}^{K}\tilde{a}_{k}=k$ exactly.
Subtracting the mean of $\tilde{a}$ introduces a uniform shift of magnitude $C=\mathcal{O}\bigl{(}\tau+L\,\lvert\lambda_{l}-\lambda^{\star}\rvert\bigr{)}$. Hence each coordinate error can be bounded as
$\lvert\tilde{a}_{k}-a_{k}^{\star}\rvert\;\leq\;\underbrace{\lvert(f^{\prime})^{-1}(-\lambda_{l}-c_{k})-(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})\rvert}_{\leq L\,\lvert\lambda_{l}-\lambda^{\star}\rvert}\;+\;\underbrace{\lvert\tilde{y}_{i}-(f^{\prime})^{-1}(-\lambda_{l}-c_{k})\rvert}_{\leq\tau}\;+\;C.$

### v2

@@ A.2 Proofs and Auxiliary Results for Section 3
    $f(x):[0,1]\rightarrow\mathbb{R}$ is defined through $f(x)=x\ln x-x$ for $x\in(0,1]$ and $f(0)=0$, which is continuous on $[0,1]$ and differentiable on $(0,1]$. Note that although $\ln x$ is not defined at $x=0$, we have $\lim_{x\rightarrow 0^{+}}x\ln x=0$. Observe that its first derivative for $x\in(0,1]$ is $\ln x$. It then follows that $(f^{\prime})^{-1}(-\lambda-c_{k})=e^{-\lambda-c_{k}}$ for all $k\in[K]$. In addition, note that $z=f^{\prime}(x)=\ln x$ ranges over the interval $z\in(-\infty,0]$. On that interval, the derivative of the inverse $\partial e^{z}/\partial z=e^{z}$ satisfies $e^{z}\leq e^{0}=1$ for all $z\in(-\infty,0]$. By the mean-value theorem, for any $z_{1},z_{2}\in[-\infty,0]$ there exists $c$ between them such that $\left|e^{z_{1}}-e^{z_{2}}\right|=e^{c}|z_{1}-z_{2}|\leq 1|z_{1}-z_{2}|$. Hence, the function $(f^{\prime})^{-1}(-\lambda-c_{k})=e^{-\lambda-c_{k}}$ is $1$-Lipschitz.
@@
    $f(x)=-\sqrt{x}$ on $(0,1]$.
    Similar to previous calculations, we have $x=(f^{\prime})^{-1}(z)=1/(4z^{2})$. Note that it must hold $z<0$ for the expression to be defined properly. Hence, with $z=-\lambda-c_{k}<0$, for all $k\in[K]$, we have $(f^{\prime})^{-1}(-\lambda-c_{k})=1/(4(-\lambda-c_{k})^{2})$. As $x\in(0,1]$ we have $z=-1/(2\sqrt{x})\in(-\infty,-1/2]$. On that interval, the derivative of the inverse is:
@@
    On the interval $z\in\left(-\infty,-\frac{1}{2}\right]$, the largest magnitude of the inverse of the derivative map $|(f^{\prime})^{-1}(z)|$ occurs at the smallest $|z|$, namely at $|z|=1/2$. This follows from the fact that $|(f^{\prime})^{-1}(z)|$ is decreasing as $|z|$ grows. Then, $\sup|(f^{\prime})^{-1}(z)|=4$, on $z\in(-\infty,-1/2]$. So, $|(f^{\prime})^{-1}(z_{1})-(f^{\prime})^{-1}(z_{2})|\leq 4|z_{1}-z_{2}|$, for all $z_{1},z_{2}\in(-\infty,-1/2]$ Therefore, $(f^{\prime})^{-1}$ is $4$-Lipschitz.
@@
$f^{\prime}(m/K)\leq-\underline{\lambda}-c_{k}\quad\forall k\in[K].$ (11)
As $f$ is strictly convex, its first derivative $f^{\prime}$ is strictly increasing, which in turn implies that its inverse $(f^{\prime})^{-1}$ is also strictly increasing. Applying the inverse function to both sides of equation 11 yields
$(f^{\prime})^{-1}(-\underline{\lambda}-c_{k})\geq(f^{\prime})^{-1}\left(f^{\prime}\left(\frac{m}{K}\right)\right)=\frac{m}{K}\quad\forall k\in[K].$
Summing over $k=1,\dots,K$ gives $\sum_{k=1}^{K}(f^{\prime})^{-1}(-\underline{\lambda}-c_{k})\geq m$. A similar argument applied to $\bar{\lambda}$ gives $\sum_{k=1}^{K}(f^{\prime})^{-1}(-\bar{\lambda}-c_{k})\leq m$. Thus, by the intermediate value theorem, there exists at least one $\lambda^{*}\in[\underline{\lambda},\bar{\lambda}]$ such that
$\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})=m.$
@@
where $\Delta=m-\sum_{k=1}^{K}(f^{\prime})^{-1}(-c_{k}-\lambda^{\star})$. For each coordinate $k$, the error induced by the difference between $\lambda$ and $\lambda^{\star}$ is dominated by the Lipschitz property equation 12 through
@@
which in turn implies that $(f^{\prime})^{-1}(-\lambda-c_{k})\geq(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})-\varepsilon/(2\sqrt{K})$. Summing over $k=1,\dots,K$ and by defining $\Delta_{\lambda}=m-\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda-c_{k})$ yields:
$\Delta_{\lambda}\leq m-\left[\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})-\frac{\varepsilon\sqrt{K}}{2}\right].$
By the definition of $\Delta$ we can bound $\Delta_{\lambda}$ as $\Delta_{\lambda}\leq\Delta+\varepsilon\sqrt{K}/2$. In the worst-case scenario when the computed error $\Delta_{\lambda}$ is close to 0, the inequality above reduces to $\Delta\geq\Delta_{\lambda}-\varepsilon\sqrt{K}/2$. Since $\Delta_{\lambda}$ is nearly 0 or in the worst-case non-positive, this yields the bound $\Delta\geq-\varepsilon\sqrt{K}/2$. Therefore, from equation 13, the error in each coordinate satisfies $|a_{k}-a_{k}^{\star}|\leq\varepsilon/\sqrt{K}$, and therefore $\|a-a^{\star}\|_{2}\leq\varepsilon$. Thus, the claim follows. ∎
@@
Define the exact residual $g(\lambda)=m-\sum_{k=1}^{K}(f^{\prime})^{-1}(-\lambda-c_{k})$ and the approximate residual $\tilde{g}(\lambda)=m-\sum_{k=1}^{K}\tilde{y}_{k}$, where each $\tilde{y}_{k}$ satisfies $|\tilde{y}_{k}-(f^{\prime})^{-1}(-\lambda-c_{k})|\leq\tau$. To measure the oracle’s perturbation, set $E(\lambda)=\tilde{g}(\lambda)-g(\lambda)=-\sum_{k=1}^{K}\left[\tilde{y}_{k}-(f^{\prime})^{-1}(-\lambda-c_{k})\right]$. Since this is a sum of $K$ individual errors, the triangle inequality gives $|E(\lambda)|\leq\sum_{k=1}^{K}|\tilde{y}_{k}-(f^{\prime})^{-1}(-\lambda-c_{k})|\leq K\tau$.
@@
Provided that $\min\{g(\underline{\lambda}),\,-g(\bar{\lambda})\}>K\tau$, the sign test on $\tilde{g}$ still selects the correct subinterval; hence the bisection converges with each evaluation of $g$ perturbed by at most $\pm K\tau$. After $l$ iterations, the midpoint $\lambda_{l}$ satisfies $\lvert\lambda_{l}-\lambda^{\star}\rvert\;\leq\;\Delta_{0}/2^{l+1}$, with $\Delta_{0}=\bar{\lambda}-\underline{\lambda}$. Upon termination at $\lambda_{l}$, define $\tilde{a}_{k}=\tilde{y}_{k}(\lambda_{l})$ for all $k\in[K]$. Finally, subtract the mean of $\tilde{a}$ to ensure that $\sum_{k=1}^{K}\tilde{a}_{k}=m$ exactly.
[…]
$\lvert\tilde{a}_{k}-a_{k}^{\star}\rvert\;\leq\;\underbrace{\lvert(f^{\prime})^{-1}(-\lambda_{l}-c_{k})-(f^{\prime})^{-1}(-\lambda^{\star}-c_{k})\rvert}_{\leq L\,\lvert\lambda_{l}-\lambda^{\star}\rvert}\;+\;\underbrace{\lvert\tilde{y}_{k}-(f^{\prime})^{-1}(-\lambda_{l}-c_{k})\rvert}_{\leq\tau}\;+\;C.$
@@
Appendix B Additional Numerical Experiments
In Sec. 3.1 we reported the per-iteration runtime of Algorithm 3. To further validate both the approach and our implementation, we plot cumulative-regret trajectories for OSMD (Algorithm 2), comparing (a) our Bisection-based projection with (b) a direct solution of the projection step via MOSEK. [^4] All code to reproduce the figures is available at https://github.com/RAO-EPFL/BOBW-CCB.
Mean cumulative regrets for the stochastic and adversarial settings are shown in Figs. 2, 3, 4, and 5, respectively, under the Tsallis and negative Shannon-entropy regularizers. We consider $m\in\{3,5\}$, $K\in\{20,40\}$ base arms, and a horizon of $T=10^{4}$ rounds.
Across all configurations, the two implementations yield indistinguishable regret curves (up to numerical tolerance), supporting the correctness of our implementation. For completeness, Tables 2 and 3 report final cumulative regret for all projection subroutines considered in the paper (Bisection, Newton, MOSEK) under the Tsallis and negative Shannon-entropy regularizers, respectively.
(a) Bisection (Ours).
(b) MOSEK.
Figure 2: Stochastic setting for $\Delta=0.0625$ (Tsallis regularizer).
(a) Bisection (Ours).
(b) MOSEK.
Figure 3: Stochastic setting for $\Delta=0.0625$ (Negative Shannon entropy regularizer).
(a) Bisection (Ours).
(b) MOSEK.
Figure 4: Adversarial setting for $\Delta=0.0625$ (Tsallis regularizer).
(a) Bisection (Ours).
(b) MOSEK.
Figure 5: Adversarial setting for $\Delta=0.0625$ (Negative Shannon entropy regularizer).
Table 2: Cumulative regret at $T=10^{4}$ for $\Delta=0.0625$ (Tsallis regularizer).
$K$ $m$ Method Stochastic $(\mu\pm\sigma)$ Adversarial $(\mu\pm\sigma)$
2020 33 Bisection (Ours) 711.3± 146.3711.3\ $\pm\$146.3 694.6± 161.2694.6\ $\pm\$161.2
Newton 726.2± 170.1726.2\ $\pm\$170.1 710.8± 130.4710.8\ $\pm\$130.4
MOSEK 734.5± 143.0734.5\ $\pm\$143.0 686.6± 88.1686.6\ $\pm\$88.1
55 Bisection (Ours) 775.7± 170.2775.7\ $\pm\$170.2 736.5± 119.7736.5\ $\pm\$119.7
Newton 761.2± 164.3761.2\ $\pm\$164.3 734.7± 121.3734.7\ $\pm\$121.3
MOSEK 758.8± 170.1758.8\ $\pm\$170.1 721.5± 137.4721.5\ $\pm\$137.4
4040 33 Bisection (Ours) 1059.4± 190.21059.4\ $\pm\$190.2 1099.7± 143.41099.7\ $\pm\$143.4
Newton 1114.8± 181.41114.8\ $\pm\$181.4 1117.4± 171.51117.4\ $\pm\$171.5
MOSEK 1095.2± 176.11095.2\ $\pm\$176.1 1101.4± 119.71101.4\ $\pm\$119.7
55 Bisection (Ours) 1303.2± 167.31303.2\ $\pm\$167.3 1312.6± 199.01312.6\ $\pm\$199.0
Newton 1314.9± 210.21314.9\ $\pm\$210.2 1309.8± 187.61309.8\ $\pm\$187.6
MOSEK 1338.0± 197.41338.0\ $\pm\$197.4 1285.6± 164.91285.6\ $\pm\$164.9
Table 3: Cumulative regret at $T=10^{4}$ for $\Delta=0.0625$ (Negative Shannon entropy regularizer).
$K$ $m$ Method Stochastic $(\mu\pm\sigma)$ Adversarial $(\mu\pm\sigma)$
2020 33 Bisection (Ours) 845.1± 160.9845.1\ $\pm\$160.9 871.5± 107.2871.5\ $\pm\$107.2
Newton 835.6± 136.9835.6\ $\pm\$136.9 874.0± 102.8874.0\ $\pm\$102.8
MOSEK 847.4± 115.7847.4\ $\pm\$115.7 887.6± 131.4887.6\ $\pm\$131.4
55 Bisection (Ours) 910.2± 164.4910.2\ $\pm\$164.4 914.8± 126.2914.8\ $\pm\$126.2
Newton 914.4± 152.4914.4\ $\pm\$152.4 900.5± 128.8900.5\ $\pm\$128.8
MOSEK 925.1± 160.6925.1\ $\pm\$160.6 893.7± 129.9893.7\ $\pm\$129.9
4040 33 Bisection (Ours) 1300.8± 149.41300.8\ $\pm\$149.4 1339.2± 115.31339.2\ $\pm\$115.3
Newton 1281.1± 118.81281.1\ $\pm\$118.8 1342.8± 101.91342.8\ $\pm\$101.9
MOSEK 1340.1± 144.51340.1\ $\pm\$144.5 1325.0± 123.51325.0\ $\pm\$123.5
55 Bisection (Ours) 1636.6± 193.41636.6\ $\pm\$193.4 1648.3± 156.41648.3\ $\pm\$156.4
Newton 1568.7± 168.51568.7\ $\pm\$168.5 1629.0± 141.81629.0\ $\pm\$141.8
MOSEK 1612.7± 168.01612.7\ $\pm\$168.0 1642.0± 170.91642.0\ $\pm\$170.9